# Kolmogorov-size particles in homogeneous and isotropic turbulence

### v1

@@ 6 Conclusion
We have investigated by direct numerical simulations the fluid-solid interaction of suspensions of Kolmogorov-size spherical particles moving in homogeneous isotropic turbulence. The work is based on both interface-resolved (PR-DNS) and one-way-coupled point-particle (PP-DNS) direct numerical simulations. In PR-DNS the presence of the particles is dealt with the immersed boundary method introduced by Hori et al. (2022). In PP-DNS the motion of the particles is described by solving the complete Maxey-Riley-Gatignol equation (Maxey & Riley, 1983; Gatignol, 1983), including the time-history Basset term and the Faxén correction. The objective of the study is twofold. On one side, we aim to shed light on how Kolmogorov-size particles influence the organisation of the velocity fluctuations. Few work have indeed considered particles with $D_{p}/\eta\approx 1$ due to the intrinsic complexity of the problem: in experiments it requires the access to sub-Kolmogorov measurements, and in simulations it requires an extremely fine grid with a resulting prohibitive computational cost. On the other side, we aim to assess the limits of the one-way-coupled PP-DNS that, despite the large number of works present in literature, have not been completely addressed yet. For this reason, we consider a portion of the parameter space that is on the edge of the range of validity of the one-way coupled point-particle models (see Brandt & Coletti, 2022). The micro-scale Reynolds number is $Re_{\lambda}\approx 140$, being large enough to ensure a proper separation of scales. The volume fraction of the suspension has been set to the small values of $\Phi_{V}=10^{-5}$ and $\Phi_{V}=10^{-3}$ to guarantee that the backreaction of the solid phase on the carrier flow is low. Two solid-to-fluid density ratios are considered, i.e. $\rho_{p}/\rho_{f}=5$ and $\rho_{p}/\rho_{f}=100$, to investigate the role of inertia.
@@
The PR-DNS shows that at the present parameters the modulation of the flow is rather low and mainly involves the smallest scales. The modulated energy spectrum $\mathcal{E}(\kappa)$ shows a multi-scaling behaviour: the classical $-5/3$ scaling in the inertial range of scales is indeed followed by a steeper $-4$ scaling, that resembles what has been observed by several authors in the context of bubbly flows (see Pandey et al., 2023). Accordingly, the scale-by-scale energy budget shows two different regimes: in the inertial range of scales the fluid-structure interaction term is negligible and the non linear term equals the dissipation rate, i.e. $\Pi(\kappa)\sim\epsilon$; at these scales energy is transferred from larger to smaller scales by means of the classical energy cascade described by Richardson and Kolmogorov. At small scales, where $\mathcal{E}(\kappa)\sim\kappa^{-4}$, the non linear term is negligible and the fluid-structure interaction term is balanced by the viscous dissipation, i.e. $\Pi_{fs}(\kappa)\sim D_{v}(\kappa)$; at these scales the energy injected into the flow by the particles is directly dissipated by viscosity. The small-scale topology of the flow has been investigated by inspecting the influence of the particles on the invariants of the velocity gradient tensor $A_{ij}=\partial u_{i}/\partial x_{j}$ (Meneveau, 2011). The effect of the solid phase on the eigenvalues of the strain-rate tensor shows that the presence of the particles favours axisymmetric compression rather than axisymmetric extension. The joint p.d.f of the second and third invariants of $A_{ij}$ reveals that, particles mainly enhance events with axial strain and vortex compression. Accordingly, the inspection of the joint p.d.f. of the second invariants of the symmetric and antisymmetric parts of $A_{ij}$ indicates that the presence of the particles favours dissipation events dominated by extensional events rather than rotational ones.
@@
By characterising the fluid-solid interaction of Kolmogorov-size particles in homogeneous isotropic turbulence, the present study aims to serve as a stepping stone for further investigations. A natural extension of this work is to use the present PR-DNS database to assess the limits of two-way-coupled PP-DNS, that account for the backreaction of the solid phase to the carrier flow; see for instance the models introduced by Gualtieri et al. (2015) and Vreman (2016). In addition, the present results may be used as a ground truth for studies in the spirit of Olivieri et al. (2014), that investigate the relevance of each term at the right-hand-sides of the MRG equation in predicting the different statistics of the particles. This knowledge will help guide the choice of suitable models for engineering applications. Eventually, it would be of interest to investigate whether Kolmogorov-size particles modulate the energy spectrum also in the inertial range of scales at larger volume fractions, influencing thus the $-5/3$ scaling range and the classical energy cascade as observed for larger particles (Chiarini & Rosti, 2024). Despite the computational challenge that would such a study would provide, the field would benefit from the investigation. Overall, the present results can be exploited for the development of improved point-particle models for the one- and two-way-coupling regimes.
@@ Appendix A The Basset time history force
Here, $\phi(z)=(\exp(z)-1)/z$, and for a given value of $m$ the coefficients $\left\{a_{i},t_{i}\right\}_{i=1}^{m}$ are chosen to minimize the error. For a detailed explanation, the reader is referred to van Hinsberg et al. (2011) and Casas et al. (2018). We choose $m=10$ and set the values of the $a_{i}$ and $t_{i}$ parameters to the ones proposed in the work of van Hinsberg et al. (2011). As suggested by van Hinsberg et al. (2011) the point-particle equation is written in a semi-implicit manner to guarantee numerical stability when integrating in time.

### v2

@@ 6 Conclusion
We have investigated by direct numerical simulations the fluid-solid interaction of suspensions of Kolmogorov-size spherical particles moving in homogeneous isotropic turbulence. The work is based on both interface-resolved (PR-DNS) and one-way-coupled point-particle (PP-DNS) direct numerical simulations. In PR-DNS the presence of the particles is dealt with the immersed boundary method introduced by Hori et al. (2022). In PP-DNS the motion of the particles is described by solving the complete Maxey-Riley-Gatignol equation (Maxey & Riley, 1983; Gatignol, 1983), including the time-history Basset term and the Faxén correction. The objective of the study is twofold. On one side, we aim to shed light on how Kolmogorov-size particles influence the organisation of the velocity fluctuations. Few work have indeed considered particles with $D_{p}/\eta\approx 1$ at a Reynolds number that is large enough to ensure a separation of scales due to the intrinsic complexity of the problem: in experiments it requires the access to sub-Kolmogorov measurements, and in simulations it requires an extremely fine grid with a resulting prohibitive computational cost. On the other side, we aim to assess the limits of the one-way-coupled PP-DNS that, despite the large number of works present in literature, have not been completely addressed yet. For this reason, we consider a portion of the parameter space that is on the edge of the range of validity of the one-way-coupled point-particle models (see Brandt & Coletti, 2022). The micro-scale Reynolds number is $Re_{\lambda}\approx 140$, being large enough to ensure a proper separation of scales. The volume fraction of the suspension has been set to the small values of $\Phi_{V}=10^{-5}$ and $\Phi_{V}=10^{-3}$ to guarantee that the backreaction of the solid phase on the carrier flow is low. Two solid-to-fluid density ratios are considered, i.e. $\rho_{p}/\rho_{f}=5$ and $\rho_{p}/\rho_{f}=100$, to investigate the role of inertia.
@@
The PR-DNS shows that at the present parameters the modulation of the flow is rather low and mainly involves the smallest scales. The modulated energy spectrum $\mathcal{E}(\kappa)$ shows a multi-scaling behaviour: the classical $\kappa^{-5/3}$ scaling in the inertial range of scales is indeed followed by a steeper $\kappa^{-4}$ scaling, that resembles what has been observed by several authors in the context of bubbly flows (see Pandey et al., 2023). Accordingly, the scale-by-scale energy budget shows two different regimes: in the inertial range of scales the fluid-structure interaction term is negligible and the nonlinear term equals the dissipation rate, i.e. $\Pi(\kappa)\sim\epsilon$; at these scales energy is transferred from larger to smaller scales by means of the classical energy cascade described by Richardson and Kolmogorov. At small scales, where $\mathcal{E}(\kappa)\sim\kappa^{-4}$, the nonlinear term is negligible and the fluid-structure interaction term is balanced by the viscous dissipation, i.e. $\Pi_{fs}(\kappa)\sim D_{v}(\kappa)$; at these scales the energy injected into the flow by the particles is directly dissipated by viscosity. The small-scale topology of the flow has been investigated by inspecting the influence of the particles on the invariants of the velocity gradient tensor $A_{ij}=\partial u_{i}/\partial x_{j}$ (Meneveau, 2011). The effect of the solid phase on the eigenvalues of the strain-rate tensor shows that the presence of the particles favours axisymmetric compression rather than axisymmetric extension. The joint probability density function of the second and third invariants of $A_{ij}$ reveals that particles mainly enhance events with axial strain and vortex compression. Accordingly, the inspection of the joint probability density function of the second invariants of the symmetric and antisymmetric parts of $A_{ij}$ indicates that the presence of the particles favours dissipation events dominated by extensional events rather than rotational ones. Our findings show that the modulation of homogeneous isotropic turbulence by Kolmogorov-size particles largely differs from what has been observed in previous studies with larger particles.
@@
By characterising the fluid-solid interaction of Kolmogorov-size particles in homogeneous isotropic turbulence, the present study aims to serve as a stepping stone for further investigations. A natural extension of this work is to use the present PR-DNS database to assess the limits of two-way-coupled and four-way-coupled PP-DNS that account for the backreaction of the solid phase to the carrier flow (see for instance the models introduced by Gualtieri et al., 2015; Vreman, 2016b), and possibly the particle-particle hydrodynamic interaction. In addition, the present results may be used as a ground truth for studies in the spirit of Olivieri et al. (2014), that investigate the relevance of each term at the right-hand-sides of the MRG equation in predicting the different statistics of the particles. This knowledge will help guide the choice of suitable models for engineering applications. Eventually, it would be of interest to investigate whether Kolmogorov-size particles modulate the energy spectrum also in the inertial range of scales at larger volume fractions, influencing thus the $\kappa^{-5/3}$ scaling range and the classical energy cascade as observed for larger particles (Chiarini & Rosti, 2024). Despite the computational challenges such a study would present, the field would benefit from the investigation. Overall, the present results can be exploited for the development of improved point-particle models for the one- and two-way-coupling regimes.
@@ Appendix A The Basset time history force
Here, $\phi(z)=(\exp(z)-1)/z$, and for a given value of $m$ the coefficients $\left\{a_{i},t_{i}\right\}_{i=1}^{m}$ are chosen to minimise the error. For a detailed explanation, the reader is referred to van Hinsberg et al. (2011) and Casas et al. (2018). We choose $m=10$ and set the values of the $a_{i}$ and $t_{i}$ parameters to the ones proposed in the work of van Hinsberg et al. (2011). As suggested by van Hinsberg et al. (2011) the point-particle equation is written in a semi-implicit manner to guarantee numerical stability when integrating in time.